OpenStudy (anonymous):

−4(x + 1) − 3 = −3(x − 4)

OpenStudy (anonymous):

distribute -4x-4-3=-3x+12 Now solve -4x-7=-3x+12 Add 3x to both sides -x-7=12 +7 +7 -x=19 x can't be a negitive so change the signs x=-19 understand?